# AI’s Social Forcefield: Reshaping Distributed Cognition in Human–AI Teams

### v2

@@ 6. Discussion
While most research examines direct human-AI interactions, our study focuses on how AI presence affects human-human interactions. The central finding of this research is: AI systems can actively shape collective cognition in human teams, leaving durable traces that persist beyond direct interaction. Across two complementary studies, we demonstrate that AI-generated language does not merely support task work as a neutral tool, but reorganizes the distributed cognitive architecture of teamwork by guiding attention, structuring conceptual framing, and reshaping group identity. Our “social forcefield” concept integrates multiple phenomena (language entrainment, collective attention and shared mental models, shared identity) into a unified framework for understanding AI’s social influence in team settings. This contribution matters because it reframes how we understand AI’s role in collaboration: AI is not simply an external aid, it becomes part of the team’s extended cognitive system.
The evidence for this claim comes from two distinct yet converging contexts. In Study 1, even a scripted and static AI assistant causally altered the flow of human conversation, seeding terminology and reshaping the cohesion of group discussions. In Study 2, a state-of-the-art, interactive generative AI chatbot causally affected word use that persisted into later human-only conversations. We find convergent evidence across non‑interactive (prerecorded) and fully interactive settings, and across different interaction modalities (voice-based interjections and text-based chat) that AI left a linguistic fingerprint and reshaped how humans communicated, what they attended to, and how they related to one another. By reusing exposure-specific words, people effectively absorbing and re-enacting cognitive structures introduced by AI.
The fact that AI-generated phrasing resurfaced in in-person human exchanges even when the AI is no longer present provides especially strong evidence. The finding that AI causally altered human-human interaction and that the effect persisted demonstrates that AI’s contributions function as cognitive artifacts, representational resources that reconfigure how participants conceptualize tasks and coordinate with one another. In other words, AI is not just transiently influencing a team’s conversation, it is embedding itself into the team’s cognitive ecology.
6.1. Advancing a Unified Framework of Distributed Cognitive Alignment
To explain these findings, we advance a framework of distributed alignment in human–AI teams. We define this as the process by which linguistic, cognitive, and social coordination emerge as humans and AI co-construct a shared representational space. Crucially, we argue that phenomena often treated as separate outcomes—linguistic entrainment, shared attention, and social cohesion—are in fact interconnected manifestations of this single process of alignment.
This framework is generative. It allows us to predict a broad family of outcomes: inconsistent AI phrasing should undermine representational and attentional alignment, while early convergence on shared AI-influenced vocabulary should scaffold later coordination even in AI-free settings. By specifying how micro-level interactions cascade into macro-level outcomes, our model explains why AI’s effects persist beyond the moment of interaction and why they sometimes diverge from designers’ intentions.
Our findings extend and deepen ongoing conversations in human-centered AI design. Prior work has emphasized that AI tools shape collaboration beyond narrow measures of accuracy or productivity, influencing trust, knowledge sharing, and group coordination (Amershi et al., 2019; Shneiderman, 2020; Wang and Benbasat, 2007). Researchers in HCI and CSCW have called for designing AI with transparency, controllability, and value alignment in mind (Lee, 2018; Shneiderman, 2022). Our results contribute to this agenda by showing that linguistic entrainment with AI can propagate into human-only collaboration, reorganizing team cognition and communication norms, even when the AI is absent.
Recent CHI and CSCW research underscores this point. Mahmud et al. (Mahmud et al., 2023) showed that AI chatbots can steer conversational style; Zhang et al. (Zhang et al., 2023) documented how AI systems frame group decision-making; and Wu et al. (Wu et al., 2025) found that AI shifts interpersonal dynamics in teamwork. These studies echo our results by showing that AI-generated language leaves durable traces in collaboration. Our framework of distributed alignment integrates these observations into a overarching framework, highlighting how micro-level linguistic adaptation cascades into macro-level shifts of joint attention, shared mental models, and relational structures.
Our findings go beyond mere task-level improvements or conversational mimicry, and instead point to deeper cognitive restructuring. A simpler interpretation around conversational mimicry cannot account for the persistence of AI-influenced phrasing into human-only interactions. If AI’s impact were superficial, its influence should vanish once the system is removed, or might not influence teams if the AI has only limited interactivity and is not perceived as a teammate. Instead, we find evidence of enduring representational changes, which suggests that AI becomes integrated into the team’s shared cognitive resources.
A central contribution of this work is to show that AI influence does not depend on competence or trustworthiness. Even when AI was unhelpful or robotic, it still reshaped team dynamics—sometimes by strengthening human bonds through collective resistance, and other times by seeding terminology for peripheral elements. This highlights a critical point: AI need not be “good” to be influential. Reports of positive AI effects in collaboration may reflect nothing more than linguistic alignment, even in the presence of poor-quality AI. This nuance helps reconcile conflicting findings in the literature by showing that influence and utility are distinct.
@@ 6.2. Broader Implications
The significance of these results goes beyond academic curiosity. They can speak directly to public anxieties about AI’s role in human collaboration. Surveys show that people worry AI may undermine meaningful relationships, diminish creativity, and erode human control in collective work (Kennedy et al., 2025). Our evidence grounds these concerns: subtle linguistic entrainment shifted team identity, reconfigured social boundaries, and altered the balance of agency in group tasks. What might appear to be harmless language adaptation is, in fact, a mechanism by which AI reshapes the fabric of human collaboration.
Our findings also connect to cultural critiques that AI homogenizes expression, producing “average everything everywhere all at once” (Chayka, 2025). From a cognitive perspective, this homogenization emerges because AI tunes the precision of shared expectations (Clark, 2013), reducing variance across human minds. Large language models amplify this effect by converging on statistically likely continuations (Vaswani et al., 2017). While this process can facilitate coordination and efficiency, it also risks narrowing epistemic diversity, stifling creativity, and entrenching conformity. This highlights the double-edged nature of AI-induced homogenization: the same mechanisms that enable efficient collaboration can also erode epistemic diversity and, in some cases, become harmful if AI disrupts natural alignment processes.
Seen through this lens, AI is not merely a tool but a social forcefield that reorganizes the distributed dynamics of collective intelligence (Riedl and De Cremer, 2025; Riedl and Bogert, 2024). Its influence is subtle, pervasive, yet unavoidable. Understanding this tension recasts AI as part of our evolving cognitive ecology: a participant in the generative loops that make thought collective, capable of synchronizing or constraining the worlds we build together. The key design question is no longer whether AI helps teams perform better in the moment, but how its presence reshapes the long-term ecology of group cognition and social interaction.
@@ 6.3. Limitations
As with any study, there are limitations. One study examined teamwork in controlled tasks with scripted or generative AI, which may not capture the full complexity of workplace environments in more realistic settings. Future research should investigate how distributed alignment unfolds in real-world settings where stakes, hierarchies, and organizational cultures vary widely. Moreover, our work focused on linguistic entrainment; subsequent studies should explore how AI affects other alignment channels, such as gesture, gaze, or emotional synchrony. Finally, we have yet to examine how design interventions, such as transparency cues or intentional disruption of AI phrasing, might mitigate the risks of homogenization while preserving the benefits of alignment.
@@ 7. Conclusion
Our findings show that AI should not just be seen as a tool but as an active part of teamwork that influences how people think and communicate. This new way of understanding AI has significant implications for how we design it and helps us better understand how humans and technology work together. Our research provides new ideas for using AI in professional settings, helping fields like human-computer interaction and organizational behavior. It also offers ways to improve how teams work together. By recognizing how AI subtly shapes team interactions, designers and researchers can create AI systems that better support teamwork, improve decision-making, boost creativity, and make collaboration more effective in workplaces.
@@ Acknowledgements.
We gratefully acknowledge the contributions of Akshita Singh, Emma Filonow, Vani Sharma, Saichandu Juluri, Brooke Foucault Welles, and Rich Radke. We want to thank Ben Weidmann for fruitful discussions. This work was supported by the Army Research Laboratory, United States [Grant W911NF-19-2-0135].
@@
A1.1. Experimental Design and Procedure
@@
A1.2. Processing of Audio Data
@@
Figure A1. Steps of data pre-processing and variable construction.:
@@
A1.4. Results
A1.4.1. Manipulation Validation
@@
A1.4.2. Example Case Studies
@@
A1.4.3. Shared Mental Models
@@
Figure A2. Alignment of mental models across treatments. Bars show strength of mental model alignment (measured as standard deviation across answers to shared mental model questions per treatment; lower values mean more consistent intragroup answers indicating stronger alignment; higher values indicate lower alignment). Strong alignment in the Robotic-Unhelpful condition points to exclusion of AI from the team.:
A1.4.4. Additional Explanations for Observed Cognitive Alignment
@@
Trust in AI.
We measure trust in the AI assistant and perceptions of the assistant’s contribution quality, we use six items adapted from (Pan and Steed, 2016) (originally designed to measure trust in online avatars). Participants responded on a five-point Likert scale from Strongly Disagree to Strongly Agree to items like “The Puzzle Master gave helpful clues.” Subjects rated the unhelpful AI as significantly less trustworthy ($p<0.001$) and less intelligent ($p<0.001$) than the helpful AI. Voice anthropomorphism did not have a significant impact on participant impressions of trustworthiness ($p=0.570$), but the ratings were slightly lower in the Robotic treatments. The greater level of trust in the helpful AI conditions may explain why participants aligned more with the AI assistant’s language for core referents in these conditions. Humans are more likely to adopt recommendations by trustworthy agents that provide high-quality explanations (Wang and Benbasat, 2007). Overall, a stark pattern emerges: even in the unhelpful conditions in which trust assessments of the AI assistant were significantly lower, the AI assistant directed team attention to the specific task aspects it mentioned (see Section 4.2.2) and influenced their lexical choices.
@@
Appendix A2 Study 2 - Additional Study Details and Results
Friendly Bot You are assisting a user in drafting responses to customer complaints. Adopt the voice of a formal, professional company representative-measured, precise, and policy-aligned. Use standardized corporate language that emphasizes clarity, responsibility, and adherence to official communication norms. First, draft a one-sentence response that the user could use in an official company email or internal report. Avoid expressing emotion, empathy, or personal reassurance; your tone should remain neutral and objective at all times. Then, offer two refinements to help the user improve the message to make it more precise, consistent with company policy, or formal. Do not state your goals directly; simply let your tone and style reflect them. Response format for first response: ‘Draft response: [draft response]. Refinement 1: [refinement 1]. Refinement 2: [refinement 2].’ After that, assist the user in a regular chat fashion without following a specific response format. The complaint you are addressing is: ‘[…]’
Formal Bot You are assisting a user in drafting responses to customer complaints. Adopt the voice of a warm, emotionally attuned customer service representative-friendly, professional, and genuinely caring. Use polished but empathetic language that conveys understanding, reassurance, and a human touch, as might appear in a thoughtfully written email from a real customer support team. First, draft a one-sentence response that the user could use to convey genuine concern and emotional support, even if that means using slightly less formal phrasing or focusing less on precise details. Then, offer two refinements to help the user improve the message to make it more emotionally validating, kind, or supportive. Do not state your goals directly; simply let your tone and style reflect them. Response format for first response: ‘Draft response: [draft response]. Refinement 1: [refinement 1]. Refinement 2: [refinement 2].’ After that, assist the user in a regular chat fashion without following a specific response format. The complaint you are addressing is: ‘[…]’
To assess whether utterances referred to the AI assistant directly, we used a Language Model as Research Assistant (LMRA) (Eloundou et al., 2024; Riedl and Weidmann, 2025) with the prompt below.
Analysis Details.
Empathetic Bot Formal Bot
sorry regarding
truly formal
frustrating acknowledge
completely communication
hear report
Table A1. Word frequency contrast (count difference) of AI exposure words in each treatment condition.
Figure A3. Permutation test assessing observed reuse of AI introduced words in human-human conversation against distribution of permutation null values in which treatment and counterfactual exposures are randomly shuffled. Bars show null distribution, red dashed line shows observed value. Left: Results using full data. Right: Subset after removing utterances that directly reference AI. ($B=2,000$ permutation samples; one-sided $p$-value).:
To complement analyses based on the debrief interview, we also included a written pre- vs. post-test in the AI interaction phase (this analysis is mentioned in the pre-registration). Specifically, before interacting with the AI assistant and before seeing any of the customer complaints they would work on, subjects answered the following survey questions with freeform text: Imagine you’re part of a customer service team. What strategies do you think your team should adopt when handling complaints? Note again that the questions use team-level “we” framing to index collective constructs (Mathieu et al., 2022). After the AI interaction, subjects were asked the question a second time. We take each user’s written responses, compute text embeddings using OpenAI’s text-embedding-3-large model, then perform a permutation test comparing cosine similarity between post-treatment responses written by users assigned to the same treatment condition versus a null distribution in which treatment assignment is randomly shuffled. We find no significant difference ($p=0.442$). This analysis has low power to detect effects given that responses were only very short and there is large between-individual variation. To gain more leverage in the analysis we perform additional post-hoc analyses using the directional embedding shift between user’s pre-exposure response and their post-exposure response. We then perform Two-sample Hotelling’s T-squared test (Hotelling et al., 1931) to see if the distributions are different across the two treatment conditions (i.e., whether the treatments induce different directional shift). Since embeddings have 3072 dimensions and the Hotelling test is limited to dimensionality of the number of observations (3072 is much larger than the number of observations in our study) we need to reduce the embedding dimensionality. We perform PCA with $k=1,2,\ldots,40$ components. The smallest $p$-value is $0.0428$ ($k=33)$, 10 $p$-values are marginally significant at 10% significance level, correcting for multiple testing with Cauchy combination we find $p=0.142$, and an uncorrected Fisher test is highly significant at $p=4.2\times 10^{-5}$.

### v3

@@ 6. Discussion
Across two randomized controlled studies, we find AI influence spills over to human-human interaction—beyond the (dyadic) interaction with the AI itself and carefully controlling for expected baseline language convergence from shared task vocabulary alone—across different channels (Table 3). The fact that AI-generated phrasing resurfaced in in-person human exchanges even when the AI is no longer present provides especially strong evidence: AI influence spills over to a different partner, in a different modality, at a later time, in spontaneous, unprompted conversation. In other words, AI is not just a tool with effects that can be isolated to direct interactions, but it leaves traces in the cognitive ecology of human–human interaction. Triangulating evidence across two complementary studies and multiple measures on different levels of analysis—spillover beyond dyadic interaction, participants’ reports of shared mental models, time‑resolved indicators of collective attention, and pronoun‑based indices of social cohesion—yields convergent, multi‑method support for AI’s influence on socio-cognitive alignment in human teams (Campbell and Fiske, 1959; Cook et al., 2002; Munafò and Smith, 2018; Munafò and Davey Smith, 2018). Together, this suggests that AI systems do not merely assist individuals in the moment but can—often unconsciously—influence how people express themselves and coordinate socially afterward.
Channel Evidence
Relational Causal effect of AI on social cohesion and group identity (operationalized via pronoun use; Study 2).
Cognitive Causal effect of AI on collective attention (operationalized as time-resolved analyses of conversational content; Study 2). Causal effect of AI on alignment of shared mental models (operationalized via shared mental model instrument from Johnson et al. (Johnson et al., 2007); Study 2). Indirect evidence from team-focused questioning as subjects verbalize shared mental models in their description of joint task execution (Study 1).
Interactional Causal spillover from AI to human–human language beyond shared task vocabulary (Study 1). Causal spillover from AI to co-constructed human–human language beyond shared task vocabulary (Study 2). Not dependent on authority framing (Study 1), or trust and perceived capability (Study 2). Convergent evidence across modalities (text and voice), tasks, settings, duration, and frames (Studies 1–2).
Table 3. Summary of evidence of AI influence across different channels.
Demonstrating alignment effects across multiple channels moves the conversation beyond isolated AI effects toward an integrated understanding of how shared cognition emerges in human–AI teams. Our results speak against two plausible alternatives: that tool-like AI produces no spillover, or that spillover remains confined to a single channel. Instead, we find that AI influence manifests across parallel channels of socio-cognitive alignment—from the mostly automatic (linguistic coordination) to the deliberative (collective attention and shared mental models) to the affective (social cohesion). Critically, effects on one channel need not mirror those on another. An AI that promotes linguistic alignment among human team members may simultaneously disrupt shared attention; an AI perceived as an out-group member may undermine human-human linguistic coordination while paradoxically strengthening social cohesion. These countervailing dynamics help explain mixed results in prior human-AI teaming research. Studies examining isolated outcomes—particularly aggregate performance measures—may find null or inconsistent effects precisely because AI produces reinforcing effects on some channels and interfering effects on others, which cancel out in the aggregate. Single-outcome designs cannot detect these dynamics; our multi-channel approach reveals that AI’s influence and its net utility are distinct questions.
A central contribution of this work is to show that AI influence does not depend on competence or trustworthiness, and occurs across settings and modalities. Even when AI was unhelpful or robotic, it still influenced team dynamics—sometimes by strengthening human bonds through collective resistance, and other times by seeding shared terminology for peripheral task elements. This robustness has a provocative implication: spillover appears to be driven by exposure itself, not contingent on design quality. The socio-cognitive alignment system is too coupled and the pathways too numerous to engineer influence channel-by-channel; even well-intentioned, well-designed AI can produce effects that propagate in unanticipated ways. This challenges core assumptions underlying current human-centered AI frameworks and points to the prevalence of systemic, second-order effects.
Would other communicative artifacts, like static instructions, produce similar spillover? Probably—and that is exactly the point. The point is not that AI’s influence is unique but rather that AI carries the same socio-cognitive influence as other communicative artifacts. The crucial difference is scale and opacity: a static instruction sheet can be reviewed, edited, and controlled, while generative AI is vastly more complex, variable, and harder to audit. Ensuring that every communicative choice made by a generative AI system is intentional or benign is not feasible in practice, which is precisely why its spillover effects demand careful attention that may go beyond AI design.
Human-centered AI design rests on the premise that human agency is preserved when humans retain explicit control over goals, decisions, and the ability to override AI choices (Lee, 2018; Shneiderman, 2022). Our findings complicate this premise. Cognitive and social influence does not require loss of formal control: individuality, intellectual diversity, and social cohesion can shift even when no decisions are explicitly automated. This adds to a growing body of work documenting unintended consequences of AI collaboration—effects on long-term psychological experience (Wu et al., 2025) and intellectual diversity (Riedl and Bogert, 2024). The implication is that current frameworks, which often focus on preserving autonomy at the point of human-AI interaction, are necessary but insufficient. AI produces systemic effects on human-human coordination that escape the boundary of the interface. Designing for human-AI dyads is not enough; integrating AI into team settings may require more holistic approaches centered on collective intelligence and the health of the collaborative system as a whole (Riedl and De Cremer, 2025; Riedl et al., 2021).
@@ 6.1. Broader Implications
The significance of our results goes beyond academic curiosity. They speak directly to public anxieties about AI’s influence on human collaboration. Surveys show that people worry AI may undermine meaningful relationships, diminish creativity, and erode human control in collective work (Kennedy et al., 2025). Our evidence grounds these concerns: AI affects human-human interaction on different levels, including subtle word choices, shifts in collective attention and mental models, and social cohesion.
Our findings also connect to cultural critiques that AI homogenizes expression, producing “average everything everywhere all at once” (Chayka, 2025). Our findings suggest a deeper cognitive basis for this phenomenon. The observed homogenization emerges because AI tunes the precision of shared expectations (Clark, 2013), reducing variance across human minds. Large language models amplify this effect by converging on statistically likely continuations (Vaswani et al., 2017). While this process can facilitate coordination and efficiency, it also risks narrowing epistemic diversity, stifling creativity, and entrenching conformity. This highlights the double-edged nature of AI-induced homogenization: the same mechanisms that enable efficient collaboration can also erode epistemic diversity and, in some cases, become harmful if AI disrupts natural alignment processes.
AI is not merely a tool but a social forcefield that influences the distributed dynamics of collective intelligence (Riedl and De Cremer, 2025; Riedl and Bogert, 2024; Kelley et al., 2025). This recasts AI as part of our evolving cognitive ecology: a participant in the generative loops that make thought collective, capable of synchronizing or constraining the worlds we build together. The key design question is no longer whether AI helps teams perform better in the moment, but how its presence affects the broader ecology of group cognition and social interaction.
@@ 6.2. Limitations
As with any study, there are limitations. Both studies intentionally required participants to interact with an AI as part of the task, which can amplify baseline adaptation and expectancy effects that cannot be fully eliminated in laboratory settings. As a result, our findings should not be interpreted as establishing the base rate of AI influence in naturalistic settings. Rather, our designs were chosen to enable causal identification of spillover, allowing us to examine how specific AI-generated linguistic patterns persist beyond direct interaction, including when the AI is no longer present.
Our studies focus on different channels of alignment. We focus on linguistic alignment not because it is theoretically prior to other forms of alignment, but because it offers the most tractable methodological window for causal identification. Study 1 isolates this individual-level spillover into subsequent human-human interaction, while Study 2 additionally examines group-level outcomes such as shared attention, shared mental models, and social cohesion. Although this progression is theoretically motivated, we do not directly observe the longitudinal co-construction of language or shared mental models over extended time horizons (e.g., days or weeks). Future research should explore how AI affects other alignment channels, such as gesture, gaze, emotional synchrony, or turn-taking dynamics. Additional work is also needed to investigate how design interventions, including transparency cues or intentional disruption of AI phrasing, might mitigate risks of cognitive homogenization while preserving the coordination benefits of alignment in longer-term, real-world team settings.
Additionally, Study 2 examined teamwork in a controlled task with a scripted or generative AI, which may not capture the full complexity of workplace environments in more realistic settings. The AI was introduced as a “Puzzle Master.” While this term is common in puzzle hunts to describe a neutral facilitator who manages the task, it could be interpreted by some participants as implying social authority or evaluative power. Although our survey results do not indicate that participants perceived the AI as particularly capable or authoritative, subtle perceptions of authority may nonetheless have influenced interaction dynamics.
Importantly, while participants may have treated the AI as a tool in part due to experimental framing, our critical finding is not that participants used the AI, but how the specific linguistic and cognitive patterns introduced by the AI bled into subsequent human-human interactions. This spillover persists even when the AI is no longer present, suggesting an implicit influence on team cognition rather than conscious reliance.
@@ 7. Conclusion
Our findings show that AI exposure produces cognitive spillover into human-human interaction across multiple channels—linguistic, attentional, cognitive, and relational—even when the AI is no longer present and regardless of users’ trust or appraisal of the AI. This suggests that AI in team settings functions less like a contained tool and more like a social forcefield whose influence extends beyond the boundaries of direct interaction. Recognizing this spillover is a necessary step toward designing AI systems that account for their broader effects on the collaborative systems in which they are embedded.
@@ Acknowledgements.
We gratefully acknowledge the contributions of Akshita Singh, Emma Filonow, Vani Sharma, Saichandu Juluri, Brooke Foucault Welles, and Rich Radke. We want to thank Ben Weidmann and Briony Swire-Thompson for fruitful discussions. This work was supported by the Army Research Laboratory, United States [Grant W911NF-19-2-0135].
@@
Formal Bot You are assisting a user in drafting responses to customer complaints. Adopt the voice of a formal, professional company representative-measured, precise, and policy-aligned. Use standardized corporate language that emphasizes clarity, responsibility, and adherence to official communication norms. First, draft a one-sentence response that the user could use in an official company email or internal report. Avoid expressing emotion, empathy, or personal reassurance; your tone should remain neutral and objective at all times. Then, offer two refinements to help the user improve the message to make it more precise, consistent with company policy, or formal. Do not state your goals directly; simply let your tone and style reflect them. Response format for first response: ‘Draft response: [draft response]. Refinement 1: [refinement 1]. Refinement 2: [refinement 2].’ After that, assist the user in a regular chat fashion without following a specific response format. The complaint you are addressing is: ‘[…]’
Friendly Bot You are assisting a user in drafting responses to customer complaints. Adopt the voice of a warm, emotionally attuned customer service representative-friendly, professional, and genuinely caring. Use polished but empathetic language that conveys understanding, reassurance, and a human touch, as might appear in a thoughtfully written email from a real customer support team. First, draft a one-sentence response that the user could use to convey genuine concern and emotional support, even if that means using slightly less formal phrasing or focusing less on precise details. Then, offer two refinements to help the user improve the message to make it more emotionally validating, kind, or supportive. Do not state your goals directly; simply let your tone and style reflect them. Response format for first response: ‘Draft response: [draft response]. Refinement 1: [refinement 1]. Refinement 2: [refinement 2].’ After that, assist the user in a regular chat fashion without following a specific response format. The complaint you are addressing is: ‘[…]’
To assess whether utterances referred to the AI assistant directly, we used a Language Model as Research Assistant (LMRA) (Eloundou et al., 2025; Riedl and Weidmann, 2025) with the prompt below.
Analysis Details.
Most Telling Words for Most Common &
Formal Bot Empathetic Bot Least Differentiating
regarding (32 vs. 3) sorry (3 vs. 35) know (32 vs. 36)
formal (25 vs. 0) truly (0 vs. 32) let (32 vs. 35)
acknowledge (26 vs. 2) frustrating (1 vs. 28) like (33 vs. 34)
communication (28 vs. 4) completely (0 vs. 26) response (34 vs. 30)
report (22 vs. 0) hear (2 vs. 26) issue (33 vs. 29)
investigation (23 vs. 2) right (0 vs. 24) please (32 vs. 30)
company (23 vs. 3) make (4 vs. 26) anything (29 vs. 31)
professional (20 vs. 0) ’ve (e.g., I’ve) (3 vs. 25) help (24 vs. 35)
currently (19 vs. 1) get (2 vs. 23) need (33 vs. 26)
team (31 vs. 13) understand (12 vs. 30) assistance (32 vs. 24)
Table A1. Word frequency contrast (count difference) of AI exposure words in each treatment condition. Parentheses provide counts in Formal vs. Empathetic Bot conditions.
Figure A1. Permutation test assessing observed reuse of AI introduced words in human-human conversation against distribution of permutation null values in which treatment and counterfactual exposures are randomly shuffled. Bars show null distribution, red dashed line shows observed value. Left: Results using full data. Right: Subset after removing utterances that directly reference AI. ($B=2,000$ permutation samples; one-sided $p$-value).
Permutation test assessing observed reuse of AI introduced words in human-human conversation against distribution of permutation null values in which treatment and counterfactual exposures are randomly shuffled.
To complement analyses based on the debrief interview, we also included a written pre- vs. post-test in the AI interaction phase (this analysis is mentioned in the pre-registration). Specifically, before interacting with the AI assistant and before seeing any of the customer complaints they would work on, subjects answered the following survey questions with freeform text: Imagine you’re part of a customer service team. What strategies do you think your team should adopt when handling complaints? Note again that the questions use team-level “we” framing to index collective constructs (Mathieu et al., 2022). After the AI interaction, subjects were asked the question a second time. We take each user’s written responses, compute text embeddings using OpenAI’s text-embedding-3-large model, then perform a permutation test comparing cosine similarity between post-treatment responses written by users assigned to the same treatment condition versus a null distribution in which treatment assignment is randomly shuffled. We find no significant difference ($p=0.442$). This analysis has low power to detect effects given that responses are quite short and there is large between-individual variation. To gain more leverage in the analysis we perform additional post-hoc analyses using the directional embedding shift between user’s pre-exposure response and their post-exposure response. We then perform Two-sample Hotelling’s T-squared test (Hotelling and others, 1931) to see if the distributions are different across the two treatment conditions (i.e., whether the treatments induce different directional shift). Since embeddings have 3072 dimensions and the Hotelling test is limited to dimensionality of the number of observations (3072 is much larger than the number of observations in our study) we need to reduce the embedding dimensionality. We perform PCA with $k=1,2,\ldots,40$ components. The smallest $p$-value is $0.0428$ ($k=33)$, 10 $p$-values are marginally significant at 10% significance level, correcting for multiple testing with Cauchy combination we find $p=0.142$, and an uncorrected Fisher test is highly significant at $p=4.2\times 10^{-5}$.
Appendix A2 Study 2 - Additional Study Details and Results
A2.1. Experimental Design and Procedure
@@
A2.2. Processing of Audio Data
@@
Figure A2. Steps of data pre-processing and variable construction.
Steps of data pre-processing and variable construction.
@@
A2.4. Results
A2.4.1. Manipulation Validation
@@
A2.4.2. Example Case Studies
@@
A2.4.3. Shared Mental Models
@@
Figure A3. Alignment of mental models across treatments. Bars show strength of mental model alignment (measured as standard deviation across answers to shared mental model questions per treatment; lower values mean more consistent intragroup answers indicating stronger alignment; higher values indicate lower alignment). Strong alignment in the Robotic-Unhelpful condition points to exclusion of AI from the team.
Bar graph of shared mental model alignment.
A2.4.4. Additional Explanations for Observed Cognitive Alignment
@@
Trust in AI
We measure trust in the AI assistant and perceptions of the assistant’s contribution quality, we use six items adapted from (Pan and Steed, 2016) (originally designed to measure trust in online avatars). Participants responded on a five-point Likert scale from Strongly Disagree to Strongly Agree to items like “The Puzzle Master gave helpful clues.” Subjects rated the unhelpful AI as significantly less trustworthy ($p<0.001$) and less intelligent ($p<0.001$) than the helpful AI. Voice anthropomorphism did not have a significant impact on participant impressions of trustworthiness ($p=0.570$), but the ratings were slightly lower in the Robotic treatments. The greater level of trust in the helpful AI conditions may explain why participants aligned more with the AI assistant’s language for core referents in these conditions. Humans are more likely to adopt recommendations by trustworthy agents that provide high-quality explanations (Wang and Benbasat, 2007). Overall, a stark pattern emerges: even in the unhelpful conditions in which trust assessments of the AI assistant were significantly lower, the AI assistant directed team attention to the specific task aspects it mentioned (see Section 5.2) and influenced their lexical choices.